Grade 12Algebra

Solve the following equations:
2(x + 4) = 12 …............

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us divide both the side by 2,
= (2(x + 4))/2 = 12/2
=x +4 = 6
By transposing 4 from LHS to RHS it becomes -4
= x =6–4
= x =2

Thanks